it takes thiannon 30 minutes to type three-fourths of a page of her term paper. at this rate, how long will it take to type a 6-page paper?
Answer by ewatrrr(24785)


Hi,
Let x represent time it will take to type a 6-page paper
Question Sets Up the Proportion
30min/.75 page = x/6 pages
Cross Multiply to solve for x
.75x = 6*30min
x = 6*30min/.75
x = 240min or 4 hrs
CHECKING our Answer
30/.75 = 40 = 240/6 = 40